OpenStudy (anonymous):

linear Algebra: Recall that P is a projection matrix if and only if P^T=P and P^2=P. Also, recall that R is a reflection matrix if and only if R^T=R and R^2=I. Finally, recall that A is an orthogonal matrix if and only if AA^T=I=A^TA. Let P be a projection matrix. Determine whether each of the following statements is true or false. 1. If is an eigenvalue for P, then =0 or =1. 2. −1 is an eigenvalue for P. 3. det(P)=0 or det(P)=1. 4. −P is a projection matrix. 5. P^2=I 6. P is a reflection matrix.

OpenStudy (anonymous):

@joemath314159

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

I see this was posted a while ago. My bad, im in the middle of packing >.<

OpenStudy (anonymous):

I am sorry about that I think 1 and 3

OpenStudy (anonymous):

@joemath314159

OpenStudy (anonymous):

You are correct. those are the only ones that are true.

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

what about -P isnt it a projection?

OpenStudy (anonymous):

@eliassaab do u have anything to add?

OpenStudy (anonymous):

You can check if -P satisfies the condition:\[P^2 = P\]or not.

OpenStudy (anonymous):

well it does -P=(-P)^2

OpenStudy (anonymous):

\[(-P)^2 = (-P)(-P) = (-1)(-1)(P^2) = P^2 =P\ne -P\]

OpenStudy (anonymous):

oh I see so the right answers are 1 and 3

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I agree 1 and 3

OpenStudy (anonymous):

Let x be an eigenvector of a a projection matrix, and c its eigenvalue. Then:\[Px = cx\]Applying P one more time gives:\[P(Px)=P(cx)\Longrightarrow P^2x=c(Px)\]but P^2=P, and Px = cx, so we obtain:

OpenStudy (anonymous):

\[P^2x=c(Px)\Longrightarrow Px=c(cx)\Longrightarrow cx=c^2x\Longrightarrow c=c^2\]which only has the solutions 0 or 1.

OpenStudy (anonymous):

For 3, note that the det(P) is the product of its eigenvalues, So it will be a product of only 1's, only 0s, or 1's and 0s. So the det will either be 1 or 0.

OpenStudy (anonymous):

ok thanks a lot for the explanation, would mind if u take a look at my new post http://openstudy.com/study#/updates/4fdbf607e4b0f2662fd226b2

OpenStudy (anonymous):

You can also say if \( \lambda\) is an eigenvalue and X is its eigenvector, then \[ PX= \lambda X\\ P^2 X = \lambda^2 X=PX=\lambda X\\ \lambda = \lambda^2 \\ \lambda = 0 \text { or } \lambda =1 \]